OpenStudy (anonymous):

X(years since 1991) Years 2,5,7,8 y(Book sales billions$) 18,21,23,24 Write an equation in slope-intercept form to represent the data

OpenStudy (ivettef365):

first find the slope, get points (2,18) and (5,21) use slope formula

OpenStudy (anonymous):

using that y1 y2 x1 x2?

OpenStudy (ivettef365):

yes

OpenStudy (ivettef365):

y2 - y1 ------- x2 - x1

OpenStudy (anonymous):

21-18 ------ 5 -2

OpenStudy (anonymous):

is that right

OpenStudy (ivettef365):

correct, so what is the slope

OpenStudy (anonymous):

1

OpenStudy (anonymous):

whats the next step

OpenStudy (ivettef365):

ok use formula y = mx + b to find b, you have m and take any points for x,y, you can use *(2,18)

OpenStudy (anonymous):

tell me if im right so far 18=1(2)=b

OpenStudy (ivettef365):

good

OpenStudy (ivettef365):

so what is b

OpenStudy (anonymous):

18=2+b 18-2=b b=16

OpenStudy (ivettef365):

ok, so the equation you are looking for is in form of y = mx + b y = x + 16

OpenStudy (anonymous):

there we go ty

OpenStudy (ivettef365):

yw :)